Ashley won $5000 in a contest. She invested some of the money at 5% simple interest and $400 less than twice that amount at 6.5%. In one year, she earned $298 in interest. How much did she invest at each rate?
Found 3 solutions by onlinetutor365.com, ilovemath, Alan3354:
Answer by onlinetutor365.com(14)

Let x = the amount of money invested at 5% simple interest
Then, 2x – 400 invested at 6.5%
Then we have
5%x + 6.5%(2x – 400) = 298
5x + 65(2x – 400) = 29800
50x + 130x – 2600 = 29800E
180x = 32400
x = 180
and 2x -400 = 2(1800) – 400 = 3200

Answer by ilovemath(8)

I think I'm cute and clever.
I'm mistaken.

Answer by Alan3354(69443)